# Self-trapping phenomenon, multistability and chaos in open anisotropic Dicke dimer

### v1

@@ Abstract
We identify the chaotic region (depicted in Fig.4(a)) from positive average Lyapunov exponent (LE) $\bar{\Lambda}$ computed for an ensemble of phase space points Lichtenberg ; Lyapunov . Furthermore, we study the decorrelator Decorr1 ; Decorr2 ; Decorr3 ; Decorr4 ; Decorr5 ; Decorr6 dynamics of spin $D(i,t)=1-\langle{\bf s}_{i}^{a}\cdot{\bf s}_{i}^{b}\rangle$, where the initial configurations $a,b$ differ by a small perturbation, $\mathbf{s}^{b}=\mathbf{s}^{a}+\delta\mathbf{s}$, and the averaging is performed over the initial states. In the chaotic regime, the decorrelator initially grows and finally saturates below unity, whereas it remains vanishingly small for regular phase space dynamics SM . As evident from Fig.4(b) the saturation value of the decorrelator $D_{s}$ serves as an alternate measure to quantify overall chaoticity, which can efficiently identify the chaotic region in parameter space. Fig.4(a) depicts the gradual transition of the chaotic region into quasi-periodic motion, characterized by a peaked distribution of frequencies and a small value of Lyapunov Exponent (LE) SM .
@@
In the current scenario of dissipative chaos, the attractor occupies a finite region of phase space. In consequence, the ST steady states can coexist with chaos, as shown over the Bloch sphere in Fig.5(a). This remarkable coexistence phenomena can be probed dynamically, as the trajectory starting close to the zero photon number and $s_{x{\rm L}}=\pm 1,s_{z{\rm R}}=1$ converges to the ST state, while chaotic dynamics can be observed for initial $N_{\downarrow\downarrow}$ state (compared in Fig.5(b1,b2)). Note that, the degree of chaos and self-trapping phenomenon can also depend on the detuning $\delta=\omega-\omega_{0}$ SM .
@@
We thank Sudip Sinha for comments and fruitful discussions. D.M. acknowledges support from Prime Minister Research Fellowship (PMRF).

### v3

@@ Abstract
We identify the chaotic region (depicted in Fig.4(a)) from the positive average Lyapunov exponent (LE) $\bar{\Lambda}$, computed for an ensemble of phase space points Lichtenberg ; Lyapunov . Furthermore, we study the decorrelator Decorr1 ; Decorr2 ; Decorr3 ; Decorr4 ; Decorr5 ; Decorr6 ; Decorr7 dynamics of spin $D(i,t)=1-\langle{\bf s}_{i}^{a}\cdot{\bf s}_{i}^{b}\rangle$, where the initial configurations $a,b$ differ by a small perturbation, $\mathbf{s}^{b}=\mathbf{s}^{a}+\delta\mathbf{s}$, and the averaging is performed over the initial states. In the chaotic regime, the decorrelator initially grows and finally saturates below unity, whereas it remains vanishingly small for regular phase space dynamics. As evident from Fig.4(b) the saturation value of the decorrelator $D_{s}$ serves as an alternate measure to quantify overall chaoticity, which can efficiently identify the chaotic region in the parameter space. Fig.4(a) depicts the gradual transition of the chaotic region into quasi-periodic motion, characterized by a peaked distribution of frequencies and a small value of the Lyapunov Exponent (LE).
@@
In the current scenario of dissipative chaos, the attractor occupies a finite region of phase space. In consequence, the ST steady states can coexist with chaos, as shown over the Bloch sphere in Fig.5(a). This remarkable coexistence phenomena can be probed dynamically, as the trajectory initialized with zero photon number and $s_{z{\rm L}}=1,s_{x{\rm R}}=\pm 1$ converges to the ST state, while chaotic dynamics can be observed for initial $N_{\downarrow\downarrow}$ state (compared in Fig.5(b1,b2)). Note that the degree of chaos and self-trapping phenomenon can also depend on the detuning $\delta=\omega-\omega_{0}$ SM .
Quantum dynamics: To this end, we study the semiclassical dynamics using truncated Wigner approximation (TWA) considering a large number of atoms, where a stochastic noise is added due to photon loss continous_time_crystal ; Limit_cycle_Hemmerich_2 ; quantum_LC_6 ; TWA_1 ; TWA_2 ; A_M_Rey_TWA . We demonstrate the signature of bistability between SSR and ASR phases within TWA, where the dynamics is attracted to either of the SR phases depending on the initial conditions. Furthermore, the self-trapping phenomenon of photons is also verified from the TWA simulation SM . Interestingly, the self-trapped limit cycle also persists in presence of quantum fluctuations up to a certain time scale, which increases with increasing number of atoms. The onset of chaos can also be unveiled from the broad frequency spectrum of individual stochastic trajectories.
Moreover, we confirm these intriguing phenomena from the full quantum dynamics, within the stochastic wavefunction approach Q_traj_1 ; Q_traj_2 for a small number of atoms. In the bistable regime of SSR and ASR phases, the quantum trajectories converge to either of the classical steady states (as shown in Fig.2(b)). Even for small number of atoms, we observe that a fraction of quantum trajectories are attracted towards the self-trapped state, confirming its signature, which we discuss in SM .
@@
Acknowledgements.
Acknowledgments: We thank Sudip Sinha for comments and fruitful discussions. D.M. acknowledges support from Prime Minister Research Fellowship (PMRF).